OpenStudy (anonymous):

find the quadratic equation whose roots are reciprocals of the roots of the equation 3x^2-20x+17

OpenStudy (anonymous):

\[3x^2-20x+17=0\]

OpenStudy (jdoe0001):

so.. how would you factor that?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what is the reciprocal of the roots in \[3x^2 =-20x+17\]

OpenStudy (jdoe0001):

well, you'd need to factor it first

OpenStudy (jdoe0001):

the exercise assumes you've covered factoring quadratics already

OpenStudy (anonymous):

\[\frac{ 3 }{ 1 }?\]

OpenStudy (anonymous):

please give me the answer if you know it i have to do 65 more pages in the next 3 hours